OpenStudy (anonymous):

Simplify the following: x^3-1 ______ x^2-x+1

OpenStudy (anonymous):

\[\frac{x^3-1}{x^2-x+1}=\frac{(x-1)(x^2-x+1)}{x^2-x+1} =x-1\] @bjschultze87